# Closing a catenary loop: the lariat chain, the string shooter, and the heavy elastica

### v1

@@ 1 Introduction
This extended comment, including a few new results, was stimulated by recent academic attention directed at a “cute” but challenging and historied problem in string mechanics, arising from its inclusion as part of problem 13 of the 11th International Physicists’ Tournament [1], in turn inspired by a demonstration by Yeany [2]. The problem is to determine the shape equilibria of a “string shooter”, a closed loop of axially moving string with a single support, subject to gravity and drag forces, to be described in detail below. As will be discussed at length, several important aspects of this and related problems, already documented throughout the literature, have been overlooked or misunderstood, and some recently proffered numerical solutions to the problem are not physically realizable. These conclusions follow from inspection of prior analytical results [3], however, errors in the initial presentation of these results must also be corrected in order to construct solutions. We will further connect with an earlier analytical result [4].
@@ 2 Overview of the problem and its solutions
Figure 1: A string shooter in operation, and its inventor [5]. The string moves axially clockwise with speed $v$. The launch mechanism provides a force $\bf{R}$}. (Annotations added to a still from [2], flipped horizontally for comparison with other figures in the present paper.)
The analytical results of [3] can be used to generate solutions for sufficiently high drag forces. For lower drag forces, there are no solutions to the problem as posed above.
@@
The equations of an axially moving catenary, including inertial and sufficiently weak drag forces, also generically admit only curves with everywhere nonzero curvature and nonvertical orientation, approaching straightness and verticality only asymptotically at infinite length [3]. One might thus seek to patch together two separate catenary solutions— an upper and a lower branch— near vertical orientation to create a closed loop. Finite-length shapes must necessarily be nonsmooth, with jumps in angle at the patching points that can be made arbitrarily small. While this can easily provide visually appealing matches with experimental shapes, indicating that catenary physics captures the dominant balances over much of the domain, such results are not physically admissible. The error in this approach is that the tension must also suffer a jump, requiring another external force that is not provided by the physical conditions of the problem. This jump in tension blows up as the jump in angle shrinks with increasing length of the curves.
However, stronger drag forces allow for physical patching, through two different scenarios, depending on the magnitude of the drag. Three regimes, separated by two bifurcations noted but misinterpreted in [3], are
@@
The results of patching will be seen in greater detail below. The symmetry or asymmetry of the resulting shapes depends on the interrelated launch and return angles of the shot string, but it is a curious property of the catenary equations that the lengths of the upper and lower curve branches will be equal, even for asymmetric solutions [4]. A given launch angle and string length will determine a particular spatial location for the patch point, and a particular return angle.
@@ 3 A bit of history
A notable report is that of the meteorologist Aitken [6], who presents a discussion and numerous drawings of observations of a driven heavy chain either hanging over one wheel (lariat chain) or passing between two wheels (string shooter). A platform underneath is employed to interfere with the motion, this interruption at times enabling the loop to achieve greater heights by effectively transforming it into a (chain) “fountain”. This is a low drag to weight ratio system. Aitken understood that acceleration and tension are the dominant forces, and that these could be perfectly balanced for any shape in the absence of gravity. Even in the presence of gravity, shapes evolve slowly, with the time scale for relaxation greatly exceeding that for a material element to traverse features of interest through axial motion. These slow shape dynamics are the primary feature of Tuck’s “Lariat Chain” kinetic sculpture developed a century later [7], many of which serve as science museum installations. [^1] The “lariat chain” [7] is an earthly thing that lives in the presence of gravity, but serves well to imperfectly illustrate the behavior of the abstract concept of a “lariat” [8], an inertial string loop in the absence of gravity. Aitken also discusses other multi-wheel arrangements, and the phenomenon whereby a freed loop will “run” along the platform for some distance.
@@
Aitken’s analysis of his experiments benefited from prior knowledge of relevant mechanics. In his own modest assessment, “The experiments do not contain much that is new, many of the problems having previously been mathematically wrought out by Sir William Thomson and others” [6]. The dominant normal force balance between centripetal acceleration and tension was well understood in the 1850s by Routh [9], who may have picked this up from the Mathematical Tripos. Both of these contributions depend linearly on curvature, which can be divided out of the normal force balance, such that no information about shape is retained and any configuration of the axially moving body is a marginally stable equilibrium. One can alternately recognize that one of two arbitrarily shaped nonlinear wave solutions of an axially moving string is “frozen” in the lab frame. The only effect of axial motion is to positively shift the value of internal tension. This is generally true for static shape equilibria in which no drag or Coriolis forces are present. In particular, catenaries with axial flow retain their shape, with tension shifted, as noted by Routh in his treatise [9], and in contemporary analyses of telegraph-laying by Airy [10] and Thomson [11]; for a review of the literature on paying out and towing cables with additional drag forces, see [3]. For more on shape-agnosticism, tension shifts, and other interesting discussions, we refer the reader to Routh [9, Chapter XIII, Articles 594-598], Gray [12, Chapter XXII, Sections 11-13], and a short note by Healey and Papadopoulos that extends this “classical, but virtually forgotten, result” [8]. A striking consequence of the shift in tension due to axial flow is the stabilization of structures lacking bending stiffness that would otherwise collapse if under compression, in particular the catenary arch. More than a century after Aitken’s demonstrations, this phenomenon was rediscovered, explained, and experimentally demonstrated by Perkins and Mote [13] and again later by others investigating the “chain fountain” or closely related phenomena [14, 15, 16].
@@
For lighter bodies and higher velocities, which have higher drag to weight ratios, drag plays a more noticeable role, affecting both shape and tension. Slow shape dynamics and wave behavior are still interesting features, constitute the remainder of the IPT problem [1], and are on full display in Yeany’s “String Shooter/Launcher” toy and science demonstration [5, 2]. An analytical solution, accompanied by some experiments, was put forward by the astronomer Gregory in the 1940s [4].222…not long before retiring from science to pursue interests in parapsychology [17, 18]. Gregory cites no previous work— later authors will reciprocate— but indicates the following origin: “It has long been known to engineers that a belt when driven at a high speed will occasionally be seen to stand up in the air in the form of a loop… The author’s attention was drawn to this problem at the Admiralty when examining a proposal to use such a loop as a means of defence against aircraft.” He expresses interest in using this problem to measure drag forces. A similar sentiment, connected with the physics of fly fishing lines, drove the recent lariat chain experiments of Judd [19]. Gregory patches upper and lower curves together to create a loop. He realizes that at lower drag the tension is clearly not patchable, and that at higher drag, the curvature becomes singular at the patch point. He thus interprets his analytical solution as being only valid for intermediate values of drag, not recognizing the potential validity of singular high drag solutions.
@@ 4 Critique of recent literature
Most of what is needed to understand the string shooter problem is present in [3], as a special case of a broader multiparameter family of solutions that include translations of the catenary at any angle with respect to gravity. While anisotropic drag linear in velocity was considered, it was noted that for purely axially moving bodies, the solutions are more generally valid, as the drag is tangential and uniform regardless of its dependence on velocity. The solutions undergo two important bifurcations in tension and shape, whose implications for patching were given above in Section 2. The first of these, occurring at a drag to weight ratio of unity, allows the inertially-shifted tension to vanish at finite length on one side. The second of these, occurring at twice this ratio, leads to finite arc length curvature blowup. The authors of [3], including one of the authors of the present paper, made two errors. First, and most importantly, they did not realize that the first bifurcation leads to an integrable singularity, allowing patching at finite length. This led to strange results involving finite tensions despite infinite lengths, which are likely incorrect and require revisiting. [^3] Contributing to this confusion were obstacles to continuing numerical solutions through vertical orientation, and the conceptualization of the solutions in terms of heteroclinic connections in “phase portraits” for angle and curvature; above the first bifurcation, the solutions reach the angle axis orthogonally instead of asymptotically approaching it, and are thus no longer unique there, but the corresponding change in appearance of the portrait is subtle. This error in interpreting the first bifurcation carried over into an aspect of the second, where it seemed that a special nonsingular solution existed right at the second bifurcation. As stated in [3], “The existence of such a solution only under such restrictive conditions on speed is surprising, and this result will remain suspect until it is observed experimentally.” This is now resolved by the recognition that the suspect result was simply wrong. Second, they made an incorrect statement to the effect that the change in tension from infinite to zero occurs at the leading rather than the trailing side of the catenary with respect to the axial flow [3, (29)], when in fact the location of this change depends on the parameters of the problem; were the statement strictly correct, we would be unable to patch one leading upper and one trailing lower side as is necessary here to create our loops. Note that transitions at both sides are possible for the broader multiparameter family, but not for the present problem, which sits along the right edge of the diagram in [3, figure 13].
@@
Bevan and Deane [20] rediscover the solution from [3] and use it to describe the “chain fountain”, a catenary arch with tangential drag forces. They don’t realize that the drag force need not be linear. They do realize that low drag solutions cannot reach vertical orientation, which they do not need for their problem, and note the significance of the first bifurcation, but do not admit solutions above the second. Abello and co-workers [21] find what appears to be a Cartesian parameterization of the solution, which they interpret as being only valid above the first bifurcation.
@@
We conclude this section by summarizing two papers from 2019 that provide numerical solutions, and our interpretation of these in the light of previously published analytical solutions and bifurcations [4, 3] which we believe clarify some of the observations. The two groups appear to be unaware of available analytical solutions and their interpretations, including the bifurcation generating the high drag curvature singularity.
@@
Taberlet and co-workers [22] consider two paradigms, guided by experiments on heavy beaded chains and light strings draped over a pulley. For heavy bodies, they ignore drag, and further ignore the difficulties associated with vertical orientation by assuming such points sit on the pulley. They also mistakenly say that the shifted tension at the bottom of the catenary is zero, which leaves nothing to balance gravity. For light bodies they ignore gravity— the source of difficulty— so that verticality is no longer forbidden. However, they neglect to realize that the only steady shapes with drag and without gravity are straight lines [3]. Numerical solutions to the steady-state balance equations are generated using an unspecified method to “avoid the singularity”. A shape is shown for high drag, and data is presented for both moderate and high drag. They also perform unsteady numerical solutions using 1000 beads and springs, for low through high drag values. They infer the presence of the first bifurcation from the behavior of their solutions, but seem to be unaware of the second bifurcation, above which most of their results lie. It is not clear whether these simulations reach steady state, or whether an effective “bending” due to discretization, or some dynamic effect, can allow closed loop solutions at low drag. They report a numerical balance between the moments due to drag and weight forces, which is as expected if no moment is supplied at the support, and there is no net angular acceleration of the loop. The effects of drag and launch angle on shape and orientation are shown. An experimental image in their supplemental material shows a “dolphin nose” shape [23] which, along with other features visible throughout, is presumably due to bending.
@@
Daerr and co-workers [24] numerically obtain a family of static string solutions for upper and lower parts of a loop for moderate and high drag, which they patch, guided by experimental shapes. Asymptotic behavior near the vertical is considered analytically; this expansion would not be possible at low drag due to the divergence of arc length at the vertical. Although bifurcations can be inferred from these expressions, they seem unaware of them, although they note qualitative differences in shape at high drag. They employ two approaches. In one, they patch upper and lower curves at the vertical, empirically rediscovering Gregory’s result of equal lengths. In the other, they create a large jump in angle by cutting off the small region of high curvature in the top part of the experimental curves (see e.g. their figure 5c). They interpret this kink as something physical and measurable, saying the solution is obtained “on condition that both the angle and the tension are allowed to be discontinuous”, but this is not physically admissible, as it requires an injection of force from some external source not present. We interpret the “apparent angle discontinuity” data reported in their figure 5a as an arbitrary choice of these authors, rather than a physical effect. They later invoke the possibility of nonlocal drag forces as the regularizer of this feature.
@@ 5 Some misconceptions
With all the above noted, let us explicitly identify a few misconceptions and some misleading language surrounding this problem that repeatedly appear in the published scientific literature or academic discussions around the IPT: [^4] Regarding extensive discussions online on video sharing websites or other social media by professional science popularizers, companies profiting from a redesign and renaming of the original invention, or enthusiastic members of the public, suffice it to say that “someone is wrong on the internet” [25].
@@
    The body is not “lifted” by drag forces; it is supported by the force supplied at the launch point. There are no lift (normal drag) forces on a purely tangentially moving body, and the integral of a constant-magnitude tangentially-directed force around a closed loop is zero. However, as has been noted in [4, 22, 24], there is a moment about the support due to drag, which will balance the moment due to gravity. But while these moments increase if the shape “lifts” to the side, they decrease again if the shape “lifts” farther upwards. While drag can “raise” the shape somewhat, a greater effect is achieved by changing the launch angle, as will be shown below in Section 6.1.555We note that supposed visual demonstrations of the necessity of air have been performed with boundary conditions that do not capture these effects clearly, namely either close to horizontal [22] or vertical [24]. The latter is compared with a lifted-up state that does not appear to be steady; Gregory [4] points out that a vertical boundary condition requires an infinite tension. Axial motion generates both inertial forces, which change the tension, and tangential drag forces, which change the tension and the shape. The former can, for example, stabilize very tall unclosed catenaries [26] in the absence of air, or make heavy chains hover slowly [6, 7]. Both this stabilization and the drag-induced bifurcations can be misinterpreted as a sudden lifting action by the air at some threshold speed.

### v2

@@ 1 Introduction
This extended comment, including a few new results, was stimulated by recent academic attention directed at a “cute” but challenging and historied problem in string mechanics, arising from its inclusion as part of problem 13 of the 11th International Physicists’ Tournament IPT19 , in turn inspired by a demonstration by Yeany YeanyYouTube . The problem is to determine the shape equilibria of a “string shooter”, a closed loop of axially moving string with a single support, subject to gravity and drag forces, to be described in detail below. As will be discussed at length, several important aspects of this and related problems, already documented throughout the literature, have been overlooked or misunderstood, and some recently proffered numerical solutions to the problem are not physically realizable. These conclusions follow from inspection of prior analytical results ChakrabartiHanna16 , however, errors in the initial presentation of these results must also be corrected in order to construct solutions. We will further connect with an earlier analytical result Gregory49 .
@@ 2 Overview of the problem and its solutions
Figure 1: A string shooter in operation, and its inventor StringLauncher . The string moves axially clockwise with speed $v$. The launch mechanism provides a force $\bf{R}$}. (Annotations added to a still from YeanyYouTube , flipped horizontally for comparison with other figures in the present paper.)
The analytical results of ChakrabartiHanna16 can be used to generate solutions for sufficiently high drag forces. For lower drag forces, there are no solutions to the problem as posed above.
@@
The equations of an axially moving catenary, including inertial and sufficiently weak drag forces, also generically admit only curves with everywhere nonzero curvature and nonvertical orientation, approaching straightness and verticality only asymptotically at infinite length ChakrabartiHanna16 . One might thus seek to patch together two separate catenary solutions— an upper and a lower branch— near vertical orientation to create a closed loop. Finite-length shapes must necessarily be nonsmooth, with jumps in angle at the patching points that can be made arbitrarily small. While this can easily provide visually appealing matches with experimental shapes, indicating that catenary physics captures the dominant balances over much of the domain, such results are not physically admissible. The error in this approach is that the tension must also suffer a jump, requiring another external force that is not provided by the physical conditions of the problem. This jump in tension blows up as the jump in angle shrinks with increasing length of the curves.
However, stronger drag forces allow for physical patching, through two different scenarios, depending on the magnitude of the drag. Three regimes, separated by two bifurcations noted but misinterpreted in ChakrabartiHanna16 , are
@@
The results of patching will be seen in greater detail below. The symmetry or asymmetry of the resulting shapes depends on the interrelated launch and return angles of the shot string, but it is a curious property of the catenary equations that the lengths of the upper and lower curve branches will be equal, even for asymmetric solutions Gregory49 . A given launch angle and string length will determine a particular spatial location for the patch point, and a particular return angle.
@@ 3 A bit of history
A notable report is that of the meteorologist Aitken Aitken1878 , who presents a discussion and numerous drawings of observations of a driven heavy chain either hanging over one wheel (lariat chain) or passing between two wheels (string shooter). A platform underneath is employed to interfere with the motion, this interruption at times enabling the loop to achieve greater heights by effectively transforming it into a (chain) “fountain”. This is a low drag to weight ratio system. Aitken understood that acceleration and tension are the dominant forces, and that these could be perfectly balanced for any shape in the absence of gravity. Even in the presence of gravity, shapes evolve slowly, with the time scale for relaxation greatly exceeding that for a material element to traverse features of interest through axial motion. These slow shape dynamics are the primary feature of Tuck’s “Lariat Chain” kinetic sculpture developed a century later LariatChain , many of which serve as science museum installations. [^1] The “lariat chain” LariatChain is an earthly thing that lives in the presence of gravity, but serves well to imperfectly illustrate the behavior of the abstract concept of a “lariat” HealeyPapadopoulos90 , an inertial string loop in the absence of gravity. Aitken also discusses other multi-wheel arrangements, and the phenomenon whereby a freed loop will “run” along the platform for some distance.
@@
Aitken’s analysis of his experiments benefited from prior knowledge of relevant mechanics. In his own modest assessment, “The experiments do not contain much that is new, many of the problems having previously been mathematically wrought out by Sir William Thomson and others” Aitken1878 . The dominant normal force balance between centripetal acceleration and tension was well understood in the 1850s by Routh Routh55 , who may have picked this up from the Mathematical Tripos. Both of these contributions depend linearly on curvature, which can be divided out of the normal force balance, such that no information about shape is retained and any configuration of the axially moving body is a marginally stable equilibrium. One can alternately recognize that one of two arbitrarily shaped nonlinear wave solutions of an axially moving string is “frozen” in the lab frame. The only effect of axial motion is to positively shift the value of internal tension. This is generally true for static shape equilibria in which no drag or Coriolis forces are present. In particular, catenaries with axial flow retain their shape, with tension shifted, as noted by Routh in his treatise Routh55 , and in contemporary analyses of telegraph-laying by Airy Airy1858 and Thomson Thomson1857both ; for a review of the literature on paying out and towing cables with additional drag forces, see ChakrabartiHanna16 . For more on shape-agnosticism, tension shifts, and other interesting discussions, we refer the reader to Routh [Routh55, , Chapter XIII, Articles 594-598], Gray [Gray59, , Chapter XXII, Sections 11-13], and a short note by Healey and Papadopoulos that extends this “classical, but virtually forgotten, result” HealeyPapadopoulos90 . A striking consequence of the shift in tension due to axial flow is the stabilization of structures lacking bending stiffness that would otherwise collapse if under compression, in particular the catenary arch. More than a century after Aitken’s demonstrations, this phenomenon was rediscovered, explained, and experimentally demonstrated by Perkins and Mote PerkinsMote89 and again later by others investigating the “chain fountain” or closely related phenomena HannaKing11 , HannaSantangelo12 , Biggins14 .
@@
For lighter bodies and higher velocities, which have higher drag to weight ratios, drag plays a more noticeable role, affecting both shape and tension. Slow shape dynamics and wave behavior are still interesting features, constitute the remainder of the IPT problem IPT19 , and are on full display in Yeany’s “String Shooter/Launcher” toy and science demonstration StringLauncher , YeanyYouTube . An analytical solution, accompanied by some experiments, was put forward by the astronomer Gregory in the 1940s Gregory49 .222…not long before retiring from science to pursue interests in parapsychology GregoryObit1 , GregoryObit2 . Gregory cites no previous work— later authors will reciprocate— but indicates the following origin: “It has long been known to engineers that a belt when driven at a high speed will occasionally be seen to stand up in the air in the form of a loop… The author’s attention was drawn to this problem at the Admiralty when examining a proposal to use such a loop as a means of defence against aircraft.” He expresses interest in using this problem to measure drag forces. A similar sentiment, connected with the physics of fly fishing lines, drove the recent lariat chain experiments of Judd JuddLariat . Gregory patches upper and lower curves together to create a loop. He realizes that at lower drag the tension is clearly not patchable, and that at higher drag, the curvature becomes singular at the patch point. He thus interprets his analytical solution as being only valid for intermediate values of drag, not recognizing the potential validity of singular high drag solutions.
@@ 4 Critique of recent literature
Most of what is needed to understand the string shooter problem is present in ChakrabartiHanna16 , as a special case of a broader multiparameter family of solutions that include translations of the catenary at any angle with respect to gravity. While anisotropic drag linear in velocity was considered, it was noted that for purely axially moving bodies, the solutions are more generally valid, as the drag is tangential and uniform regardless of its dependence on velocity. The solutions undergo two important bifurcations in tension and shape, whose implications for patching were given above in Section 2. The first of these, occurring at a drag to weight ratio of unity, allows the inertially-shifted tension to vanish at finite length on one side. The second of these, occurring at twice this ratio, leads to finite arc length curvature blowup. The authors of ChakrabartiHanna16 , including one of the authors of the present paper, made two errors. First, and most importantly, they did not realize that the first bifurcation leads to an integrable singularity, allowing patching at finite length. This led to strange results involving finite tensions despite infinite lengths, which are likely incorrect and require revisiting. [^3] Contributing to this confusion were obstacles to continuing numerical solutions through vertical orientation, and the conceptualization of the solutions in terms of heteroclinic connections in “phase portraits” for angle and curvature; above the first bifurcation, the solutions reach the angle axis orthogonally instead of asymptotically approaching it, and are thus no longer unique there, but the corresponding change in appearance of the portrait is subtle. This error in interpreting the first bifurcation carried over into an aspect of the second, where it seemed that a special nonsingular solution existed right at the second bifurcation. As stated in ChakrabartiHanna16 , “The existence of such a solution only under such restrictive conditions on speed is surprising, and this result will remain suspect until it is observed experimentally.” This is now resolved by the recognition that the suspect result was simply wrong. Second, they made an incorrect statement to the effect that the change in tension from infinite to zero occurs at the leading rather than the trailing side of the catenary with respect to the axial flow [ChakrabartiHanna16, , (29)], when in fact the location of this change depends on the parameters of the problem; were the statement strictly correct, we would be unable to patch one leading upper and one trailing lower side as is necessary here to create our loops. Note that transitions at both sides are possible for the broader multiparameter family, but not for the present problem, which sits along the right edge of the diagram in [ChakrabartiHanna16, , figure 13].
@@
Bevan and Deane BevanDeane19 rediscover the solution from ChakrabartiHanna16 and use it to describe the “chain fountain”, a catenary arch with tangential drag forces. They don’t realize that the drag force need not be linear. They do realize that low drag solutions cannot reach vertical orientation, which they do not need for their problem, and note the significance of the first bifurcation, but do not admit solutions above the second. Abello and co-workers Abello20 find what appears to be a Cartesian parameterization of the solution, which they interpret as being only valid above the first bifurcation.
@@
We conclude this section by addressing the relevant portions of two papers from 2019 that provide numerical solutions, and our interpretation of these in the light of previously published analytical solutions and bifurcations Gregory49 , ChakrabartiHanna16 which we believe clarify some of the observations. The two groups appear to be unaware of available analytical solutions and their interpretations, including the bifurcation generating the high drag curvature singularity.
@@
Taberlet and co-workers Taberlet19 consider two paradigms, guided by experiments on heavy beaded chains and light strings draped over a pulley. For heavy bodies, they ignore drag, and further ignore the difficulties associated with vertical orientation by assuming such points sit on the pulley. They mistakenly say that the shifted tension at the bottom of the catenary is zero, which leaves nothing to balance gravity. For light bodies they ignore gravity— the source of difficulty— so that verticality is no longer forbidden. However, the only steady shapes with drag and without gravity are straight lines ChakrabartiHanna16 . For general cases between these limits, numerical solutions to the steady-state balance equations are generated using an unspecified method to “avoid the singularity”. A shape is shown for high drag, and data is presented for both moderate and high drag. They also perform dynamic numerical simulations using 1000 beads and springs, for low through high drag values. They infer the presence of the first bifurcation from the behavior of their solutions, but seem to be unaware of the second bifurcation, above which most of their results lie. It is not clear whether the low drag simulations reach steady state, or whether an effective “bending” due to discretization, or some dynamic effect, can allow closed loop solutions below the first bifurcation. They report a numerical balance between the moments due to drag and weight forces, which is as expected if no moment is supplied at the support, and there is no net angular acceleration of the loop. The effects of drag and launch angle on shape and orientation are shown. An experimental image in their supplemental material shows a “dolphin nose” shape Judddolphin which, along with other features visible throughout, is presumably due to bending stiffness.
@@
Daerr and co-workers Daerr19 numerically obtain a family of static string solutions for upper and lower parts of a loop for moderate and high drag, which they patch, guided by experimental shapes. Asymptotic behavior near the vertical is considered analytically; this expansion would not be possible at low drag due to the divergence of arc length at the vertical. Although bifurcations can be inferred from these expressions, they seem unaware of them, although they note qualitative differences in shape at high drag. They employ two approaches. In one, they patch upper and lower curves at the vertical, empirically rediscovering Gregory’s result of equal lengths. In the other, they create a large jump in angle by cutting off the small region of high curvature in the top part of the experimental curves (see e.g. their figure 5c). They interpret this kink as something physical and measurable, saying the solution is obtained “on condition that both the angle and the tension are allowed to be discontinuous”, but this is not physically admissible, as it requires an injection of force from some external source not present. We interpret the “apparent angle discontinuity” data reported in their figure 5a as an arbitrary choice of these authors, rather than a physical effect. They later invoke the possibility of nonlocal drag forces as the regularizer of this feature.
@@ 5 Some misconceptions
With all the above noted, let us explicitly identify a few misconceptions and some misleading language surrounding this problem that repeatedly appear in the published scientific literature or academic discussions around the IPT: [^4] Regarding additional extensive discussions online on video sharing websites or other social media by professional science popularizers, companies profiting from a redesign and renaming of the original invention, or enthusiastic members of the public, suffice it to say that “someone is wrong on the internet” xkcd .
@@
    The body is not “lifted” by drag forces; it is supported by the force supplied at the launch point. There are no lift (normal drag) forces on a purely tangentially moving body, and the integral of a constant-magnitude tangentially-directed force around a closed loop is zero. However, as has been noted in Gregory49 , Taberlet19 , Daerr19 , there is a moment about the support due to drag, which will balance the moment due to gravity. But while these moments increase if the shape “lifts” to the side, they decrease again if the shape “lifts” farther upwards. While drag can “raise” the shape somewhat, a quite significant effect is achieved by changing the launch angle, as will be shown below in Section 6.1.555We note that visual demonstrations of the necessity of air have been performed with boundary conditions that will not capture such effects clearly, namely either close to horizontal Taberlet19 or vertical Daerr19 . The latter is compared with a lifted-up state that does not appear to be steady; Gregory Gregory49 points out that a vertical boundary condition requires an infinite tension. Axial motion generates both inertial forces, which change the tension, and tangential drag forces, which change the tension and the shape. The former can, for example, stabilize very tall unclosed catenaries MillerYouTube in the absence of air, or make heavy chains hover slowly Aitken1878 , LariatChain . Both this stabilization and the drag-induced bifurcations can be misinterpreted as a sudden lifting action by the air at some threshold speed.